# Ammann Bars for Octagonal Tilings

### v4

@@ 3.3 Defining the prototiles
Let $\Delta\in\mathcal{D}_{i}$, $x\in\mathbb{R}^{4}$ such that $\pi(x)\in\Delta$, and $\Delta^{\prime}\in\mathcal{D}_{i}$ which is closest to $\Delta$ (Fig. 9). Then the distance from $\pi(x)$ to its orthogonal projection $\pi^{\prime}(x)$ on $\Delta^{\prime}$ is at most $d_{1}$. Besides, the distance between two vertices lying on $\Delta^{\prime}$ is at most $d_{2}:=\max(||\pi(\lfloor p_{i}\rfloor)||,||\pi(\lceil p_{i}\rceil)||)$. Indeed, if $y\in\mathbb{Z}^{4}\cap(E+[0,1]^{4})$ then $y+p_{i}\in E+[0,1]^{4}$ and has three integer coordinates so that it lies on an edge of $\mathbb{Z}^{4}$ (seen as a grid in $\mathbb{R}^{4}$), between $y+\lfloor p_{i}\rfloor$ and $y+\lceil p_{i}\rceil$; now at least one of these two points is in $\mathbb{Z}^{4}\cap(E+[0,1]^{4})$, therefore its projection is also a vertex of the tiling, which lies on $\Delta^{\prime}$ (since $\pi(p_{i})$, $\pi(\lfloor p_{i}\rfloor)$ and $\pi(\lceil p_{i}\rceil)$ are collinear). Hence the distance between $\pi^{\prime}(x)$ and the closest vertex $\pi(y)$ of the tiling which lies on $\Delta^{\prime}$ is at most $d_{2}/2$. As a result, $\text{dist}(\pi(x),\pi(y))\leq d:=\sqrt{d_{1}^{2}+d_{2}^{2}/4}$, i.e at least one vertex on $\Delta^{\prime}$ is in the ball $B(\pi(x),d)$. Consequently, measuring the intervals around a line $\Delta$ in the $d$-maps of the tiling is enough to list all possible intervals between two consecutive lines in the whole tiling. Since the $d$-atlas is finite, so is the number of intervals. $\square$
@@ 4 Tiling with an FP-tileset
For instance, every tiling composed with tiles of $\mathcal{C}$ has the same subperiods as Cyrenaic tilings. If the above proof seems obscure, here is an example of what happens in the shadows. We observe that each $i$-shadow is periodic with period $q_{i}:=\omega_{i}(p_{i})$ where $p_{i}$ is the $i$-subperiod of Cyrenaic tilings. This is shown in Figure 12.
@@
In each shadow there are three original (non-decorated) tiles, each of which can appear in different versions when taking the decorations into account. For each $i$-shadow here we only look at the decorations in direction $i$, where we have the continuity of the lines (other decorations are irrelevant here). All possible tiles are given on the top row, and following the arrows from each tile one can see all different possibilities [^7] Remember that a line passes through every vertex, in each direction. to place other tiles in order to continue the line directing the red vector. For each shadow, the vector is the same for all possible configurations, which means that the shadow is periodic, and we find exactly the subperiods of Cyrenaic tilings.

### v5

@@ 3.3 Defining the prototiles
Let $\Delta\in\mathcal{D}_{i}$, $x\in\mathbb{R}^{4}$ such that $\pi(x)\in\Delta$, and $\Delta^{\prime}\in\mathcal{D}_{i}$ which is closest to $\Delta$ (Fig. 9). Then the distance from $\pi(x)$ to its orthogonal projection $\pi^{\prime}(x)$ on $\Delta^{\prime}$ is at most $d_{1}$. Next, the distance between two vertices lying on $\Delta^{\prime}$ is at most $d_{2}:=\max(||\pi(\lfloor p_{i}\rfloor)||,||\pi(\lceil p_{i}\rceil)||)$. Indeed, if $y\in\mathbb{Z}^{4}\cap(E+[0,1]^{4})$ then $y+p_{i}\in E+[0,1]^{4}$ and has three integer coordinates so that it lies on an edge of $\mathbb{Z}^{4}$ (seen as a grid in $\mathbb{R}^{4}$), between $y+\lfloor p_{i}\rfloor$ and $y+\lceil p_{i}\rceil$; now at least one of these two points is in $\mathbb{Z}^{4}\cap(E+[0,1]^{4})$, therefore its projection is also a vertex of the tiling, which lies on $\Delta^{\prime}$ (since $\pi(p_{i})$, $\pi(\lfloor p_{i}\rfloor)$ and $\pi(\lceil p_{i}\rceil)$ are collinear). Hence the distance between $\pi^{\prime}(x)$ and the closest vertex $\pi(y)$ of the tiling which lies on $\Delta^{\prime}$ is at most $d_{2}/2$. As a result, $\text{dist}(\pi(x),\pi(y))\leq d:=\sqrt{d_{1}^{2}+d_{2}^{2}/4}$, i.e at least one vertex on $\Delta^{\prime}$ is in the ball $B(\pi(x),d)$. Consequently, measuring the intervals around a line $\Delta$ in the $d$-maps of the tiling is enough to list all possible intervals between two consecutive lines in the whole tiling. Since the $d$-atlas is finite, so is the number of intervals. $\square$
@@ 4 Tiling with an FP-tileset
For instance, every tiling composed with tiles of $\mathcal{C}$ has the same subperiods as Cyrenaic tilings. If the above proof seems opaque, here is an example of what happens in the shadows. We observe that each $i$-shadow is periodic with period $q_{i}:=\omega_{i}(p_{i})$ where $p_{i}$ is the $i$-subperiod of Cyrenaic tilings. This is shown in Figure 12.
@@
In each shadow there are three original (non-decorated) tiles, each of which can appear in different versions when taking the decorations into account. For each $i$-shadow here we only look at the decorations in direction $i$, where we have the continuity of the lines (other decorations are irrelevant here). All possible tiles are given on the top row, and following the arrows from each tile one can see all of the different options [^7] Remember that a line passes through every vertex, in each direction. for placing other tiles in order to continue the line in the direction of the red vector. For each shadow, the vector is the same for all possible configurations, which means that the shadow is periodic, and we find exactly the subperiods of Cyrenaic tilings.